A uniform cube is free to turn about one edge which is horizon. Find in terms of a seconds pendulum the length of the edge , so that it may execute a complete oscillating in which secs


1
Expert's answer
2021-03-01T12:47:19-0500


As we know seconds pendulum has a period of oscillation is 2 seconds


Let L be the length of the edge

and g =Accelaration due to gravity


Time period, "T=2\pi(\sqrt{\dfrac{L}{g}})"


"\Rightarrow 2=2\pi\sqrt{\dfrac{L}{g}}"


"\Rightarrow \dfrac{L}{g}=\dfrac{1}{\pi^2}"


"\Rightarrow L=\dfrac{g}{\pi^2}"


Hence This is the Length of the edge.